OpenStudy (love10129151):

8. The cost of four scarves and six hats is $52.00. The cost of two hats is $1.00 more than the cost of one scarf. What is the cost of one scarf? (1 point) $4.00 $5.00 $6.00 $7.0

OpenStudy (anonymous):

4 scarfs, 4s 6 hats, 6h 6h + 4s =$52 dollars 2 hats, 2h 1 scarf, 1s 2h = $1 + 1s

OpenStudy (love10129151):

so the anwers is

OpenStudy (anonymous):

that answer is... what did You get when you solved for s?

OpenStudy (love10129151):

is tjhe answer 4.00

OpenStudy (anonymous):

try again, did you use substitution to solve the system of equations? I can show you how to do this if you don't know how

OpenStudy (anonymous):

there is nothing wrong with not knowing how to do something, it just means you haven't learned it.

OpenStudy (love10129151):

yes you can show me

OpenStudy (love10129151):

ya

OpenStudy (anonymous):

here are two equation we made: 6h + 4s = 52 2h =1 + 1s we want to make one of the two equations equal to h, lets do it to the second equation. so we will need to get rid of the 2 infront of the h, can you do this, or do you need help?

OpenStudy (love10129151):

so the answer is 6.00

OpenStudy (anonymous):

not quite.. do you wanna show me what you did to solve the equations so I can help you faster and just show you where the mistake was rather than explain all the steps.

OpenStudy (love10129151):

is the anwser 5.00

OpenStudy (anonymous):

i'm sorry, i don't teach this way. i'll help you if you want to try, but i'm not going to do the guessing game with you..

OpenStudy (love10129151):

im guessing it 7.00